OpenStudy (anonymous):

find dy/dx of (x+y)^3=x^3+y^3

OpenStudy (amistre64):

the process is pretty simple with the derivative rules .. what is the issue you are having?

OpenStudy (anonymous):

the answer im getting doesnt match the back of the book i got dy/dx=[3x ^{2}-3x-3y \div3x+3y-3y ^{2}\]

OpenStudy (amistre64):

can you step me thru it so I can see where you may be going astray at?

OpenStudy (amistre64):

it may just be a matter of simplification as well

OpenStudy (anonymous):

ok heres wht i did

OpenStudy (anonymous):

also if u could help me with this one, i thought i did it but once again my answer doesnt match the back of the books. \[y ^{2}=x ^{2}-49divx ^{2}+49\]

OpenStudy (amistre64):

\[(x+y)^3 = x^3+y^3\] \[3(x+y)^2~(x+y)' = 3x^2x'+3y^2y'\] \[3(x+y)^2~(x'+y') = 3x^2x'+3y^2y'\] \[x'=\frac{dx}{dx}=1\] \[3(x+y)^2~(1+y') = 3x^2+3y^2y'\] \[(x+y)^2+y'(x+y)^2 = x^2+y^2y'\] \[y'(x+y)^2-y^2y' = x^2-(x+y)^2\] \[y'[(x+y)^2-y^2] = x^2-(x+y)^2\] \[y' = \frac{x^2-(x+y)^2}{(x+y)^2-y^2}\] \[y' = \frac{[x-(x+y)]~[x+(x+y)]}{[(x+y)+y]~[(x+y)-y]}\] \[y' = \frac{[-y]~[2x+y]}{[x+2y]~[x]}\] doesnt look to be any common factors to pull out

OpenStudy (amistre64):

i think you ignored a ^2 at the start of it

OpenStudy (anonymous):

duh. yes i forgot that squared. thnk you

OpenStudy (anonymous):

wait im confused as to how all the numbers got cancled out?

OpenStudy (amistre64):

\[y ^{2}=\frac{x ^{2}-49}{x ^{2}+49}\] \[y ^{2}=(x ^{2}-49)(x ^{2}+49)^{-1}\] now its just the product rule \[2y~y'=2xx'(x ^{2}+49)^{-1}-2xx'(x ^{2}-49)(x ^{2}+49)^{-2}\] divide out the 2 and let x' = 1 \[y~y'=x(x ^{2}+49)^{-1}-x(x ^{2}-49)(x ^{2}+49)^{-2}\] we can factor out that (x^2+49)^2 if need be \[y~y'=\frac{x(x ^{2}+49)-x(x ^{2}-49)}{(x ^{2}+49)^{2}}\] \[y'=\frac{x(x ^{2}+49)-x(x ^{2}-49)}{y(x ^{2}+49)^{2}}\] \[y'=\frac{x[(x ^{2}+49)-(x ^{2}-49)]}{y(x ^{2}+49)^{2}}\] \[y'=\frac{x(49+49)}{y(x ^{2}+49)^{2}}\]

OpenStudy (amistre64):

3 is a common factor to all the terms; dividing both sides by 3 cancels them out

OpenStudy (amistre64):

\[\frac{\cancel3(x+y)^2~(x'+y') = \cancel3x^2x'+\cancel3y^2y'}{\cancel3}\] \[(x+y)^2~(x'+y') = x^2x'+y^2y'\]

OpenStudy (amistre64):

what does the book give us?

OpenStudy (anonymous):

ok i got it now. and the book gives the same answer you got

OpenStudy (amistre64):

yay!!

OpenStudy (anonymous):

now im working on the second problem i gave you could u work it through using the quotient rule, cuz thats how i did it but didnt get the right answer

OpenStudy (amistre64):

i despise the quotient rule ... its just got to many moving parts to me. a qoutient is just another form of a product, and the product rule is sooo much easier to keep track of to me

OpenStudy (anonymous):

actually i think i found where i messed up, i forgot the bottom half of the quotient rule. let me work it thru one mroe time

OpenStudy (amistre64):

\[\frac tb=tb^{-1}\] \[[\frac tb]'=\frac{bt'-b't}{b^2}\] \[[tb^{-1}]'=t'b^{-1}+tb'^{-1}\] \[[tb^{-1}]'=t'b^{-1}-tb'b^{-2}\] \[[tb^{-1}]'=\frac{1}{b^2}(t'b-tb')\] \[[tb^{-1}]'=\frac{t'b-tb'}{b^2}\]

OpenStudy (anonymous):

ok im stuck on the making of the produuct rule. i got everything up until the taking the derviative of (x^2+49)^-1

OpenStudy (anonymous):

would we use the chain rule as well?

OpenStudy (amistre64):

power rule, drop the -1, and subtract 1 chain rule pops out the derivative of the innards

OpenStudy (amistre64):